A person takes a trip, driving with a constant speed of 99.5 km/h, except for a 22.0-min rest stop. If the person's average speed is 73.4 km/h, find the following.

(a) How much time is spent on the trip?

(b) How far does the person travel?


1
Expert's answer
2021-02-09T18:33:36-0500

a) By the definition of the average speed, we have:


"v_{avg}=\dfrac{d}{t_{tot}},"

here, "v_{avg}" is the person's average speed, "d" is the total distance covered by the person, "t_{tot}" is total time spent on the trip.

Then, from this equation, we can find "t_{tot}":


"t_{tot}=\dfrac{d}{v_{avg}}."

We can find "d" from the formula:


"d=vt,"

here, "v" is the constant speed of the person.

The total time for the entire trip can be found as follows:


"t_{tot}=t+t_{rest},"

here, "t_{rest}" is the rest stop time.

Then, we get:


"t+t_{rest}=\dfrac{vt}{v_{avg}},""t=\dfrac{t_{rest}}{\dfrac{v}{v_{avg}}-1},""t=\dfrac{22\ min\cdot\dfrac{1\ h}{60\ min}}{\dfrac{99.5\ \dfrac{km}{h}}{73.4\ \dfrac{km}{h}}-1}=1.03\ h."

Finally, we can find total time spent on the trip:


"t_{tot}=t+t_{rest}=1.03\ h+22\ min\cdot\dfrac{1\ h}{60\ min}=1.4\ h."

b) Then, we can find the total distance covered by the person:


"d=vt=99.5\ \dfrac{km}{h}\cdot1.03\ h=102.5\ km."